# Representation theorems for dynamic convex risk measures

### v3

@@ 1 Introduction
where $\mathcal{Q}:=\{Q^{q}:Q^{q}$ is a probability equivalent to $P$ such that $dQ/dP=\exp\{\int_{0}^{t}q_{s}ds-\frac{1}{2}\int_{0}^{t}|q_{s}|^{2}ds\}\}.$ Using the representation theorem established in this paper (see (3.2) in Theorem 3.3) and some results of Fan et al. [13], we show that any convex DRM satisfying (1.3) admits a dual representation as in (1.6). Compared to [10], our result holds for unbounded terminal variables, and the function $f$ can be determined by the Legendre-Fenchel transformation of the generator $g$ in (3.2). Furthermore, for coherent DRMs satisfying (1.3), we obtain a dual representation as in (1.5), where the terminal variable $\xi$ is unbounded and $\mathcal{Q}$ is characterized by the subdifferential $\partial g(t,0)$. Our method relies on the representation (3.2) and therefore requires (1.3).
@@ 2 Quadratic $g$-expectations
${\cal{G}}^{\mu,\nu}_{\theta}:=\{g\in{\cal{G}}:$ $dt\times dP$-$a.e.,$ $\forall z,\tilde{z}\in\mathbf{R}^{d}$, $\forall\theta\in(0,1)$, $g(t,z)-\theta g(t,\tilde{z})\leqslant\mu|z-\tilde{z}|+\frac{\nu}{1-\theta}|z-\tilde{z}|^{2}\}$;
@@
$\leqslant\mu|z-\tilde{z}|+\frac{\nu}{1-\theta}|z-\tilde{z}|^{2}+(1-\theta)(\mu+\nu)(1+|z|^{2}+|\tilde{z}|^{2}).$
which implies that
$\lim_{\tilde{z}\rightarrow z}|g(t,z)-g(t,\tilde{z})|\leqslant(1-\theta)(\mu+\nu)(1+2|z|^{2})\rightarrow 0,\quad\text{as}\ \theta\rightarrow 1.$
@@
Let $Y_{t}$ be a $g^{\mu,\nu}$-submartingale. For any $0\leqslant s\leqslant t\leqslant T$, by Zheng et al. [39, Corollary 3.6], we have
$Y_{s}\leqslant\mathcal{E}^{\mu,\nu}[Y_{t}|{\cal{F}}_{s}]=\frac{1}{2\nu}\ln(E_{Q}[\exp(2\nu Y_{t})|{\cal{F}}_{s}]).$
@@
Let $Y_{t}$ be a $g^{-\mu,-\nu}$-supermartingale. Since for any $0\leqslant s\leqslant t\leqslant T$,
$Y_{s}\leqslant-{\cal{E}}^{-\mu,-\nu}[Y_{t}|{\cal{F}}_{s}]={\cal{E}}^{\mu,\nu}[-Y_{t}|{\cal{F}}_{s}],$
we get that $Y_{t}$ is a $g^{\mu,\nu}$-submartingale. Thus, $Y_{t}$ admits an RCLL modification. ∎
@@
Since for any $0\leqslant s\leqslant t\leqslant T$, $\exp(2\nu Y_{s})\leqslant\exp(2\nu\mathcal{E}^{\mu,\nu}[Y_{t}|{\cal{F}}_{s}])$, by (2.2), we get that $\exp(2\nu Y_{t})$ is a $g^{\mu,0}$-submartingale. Then, by Peng [29, Theorem 3.3], there exists a pair $(z_{t},a_{t})\in H_{d}^{2}\times{\cal{A}}$, such that
@@
$Y_{t}=Y_{T}+$ $\int_{t+}^{T}\frac{1}{2\nu y_{s-}}\left(\mu|z_{s}|ds-da_{t}-z_{s}\cdot dB_{s}\right)+\int_{t+}^{T}\frac{|z_{s}|^{2}}{4\nu|y_{s}|^{2}}ds-\sum_{s\in(t,T]}\left(\Delta Y_{s}-\frac{1}{2\nu y_{s-}}\Delta y_{s}\right)$
@@ A DRM is called a regular DRM, if for all $\xi\in L^{0}(\mathcal{F}_{T})$, it satisfies (r7):
$\rho^{\mathrm{ent}}(\xi)=\mathcal{E}^{0,\nu}[-\xi|\mathcal{F}_{t}],\quad t\in[0,T],$
@@
In particular, when $\xi\in L^{\infty}(\mathcal{F}_{T})$, by (Appendix: The proof of (6.9)) in the Appendix, $\mathcal{Q}_{\xi,f}$ in (3.4) can be reduced to $\mathcal{Q}:=\{Q^{q}:$ $dQ^{q}/dP=\theta^{q}_{T}:=\int_{0}^{T}q_{s}\cdot dB_{s}-\frac{1}{2}\int_{0}^{T}|q_{s}|^{2}ds$ with $q_{t}\in\mathcal{H}_{d}^{BMO}\}.$ For $Q^{q}\in\mathcal{Q}$, the conditional relative entropy of $Q^{q}$ with respect to $P$:
@@ 5 A Doob-Meyer decomposition for $\rho$-supermartingales
$g(t,z)-\theta g(t,\tilde{z})\leqslant\mu|z-\tilde{z}|+\frac{\nu}{1-\theta}|z-\tilde{z}|^{2}.$
@@
Since $\rho$ satisfies (r5) and (r6) on $\mathcal{L}^{\exp}_{T}$, by Lemma 2.3(ix), we get that $g\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}_{subl}$. This implies that for all $z\in\mathbf{R}^{d}$ and $\beta>0$, $\beta|g(z)|=|g(\beta z)|\leqslant\beta\mu|z|+\beta^{2}\nu|z|^{2}$, and thus
$|g(z)|\leqslant\mu|z|+\beta\nu|z|^{2}.$
@@
where $B^{\tilde{q}^{n}}_{t}:=B_{t}-\int_{0}^{t}{\tilde{q}^{n}}_{s}ds$ is a standard Brownian motion under the probability $Q^{\tilde{q}^{n}}$ satisfying $Q^{\tilde{q}^{n}}/dP=\theta^{\tilde{q}^{n}}_{T}$. It follows from (Appendix: The proof of (6.9)) that there exists a constant $C$ independent of $n$ such that $E[\theta^{\tilde{q}^{n}}_{T}\ln(1+\theta^{\tilde{q}^{n}}_{T})]\leqslant C.$ Then by Protter [31, Theorem 11], we get that $\{\theta^{\tilde{q}^{n}}_{T}\}_{n\geqslant 1}$ is uniformly integrable, and then by Vitali convergence theorem, we have $E[\theta^{\tilde{q}}_{T}]=1$, which implies that $\theta^{\tilde{q}}_{t}$ is a uniformly integrable martingale. Moreover, by Fatou’s lemma, we also have
@@
By (Appendix: The proof of (6.9)), (a.10), Fatou’s lemma and the fact that $f\geqslant 0$, we have
@@
From (a.10) and (a.11), it follows that $\tilde{q}_{t}\in\mathcal{Q}_{(\xi,f)}$. Then, by (Appendix: The proof of (6.9))-(a.11) and domination convergence theorem, we have
@@
Since $f(t,q_{t})\geqslant q_{t}\cdot Z_{t}-g(t,Z_{t}),$ by (Appendix: The proof of (6.9)) and similar arguments as in (Appendix: The proof of (6.9))-(a.10), we can obtain (a.13), which together with (a.12) and Definition 2.2(i) gives (6.9). ∎

### v4

@@ 1 Introduction
where $\mathcal{Q}:=\{Q^{q}:Q^{q}$ is a probability equivalent to $P$ such that $dQ/dP=\exp\{\int_{0}^{t}q_{s}ds-\frac{1}{2}\int_{0}^{t}|q_{s}|^{2}ds\}\}.$ Using the representation theorem established in this paper (see (3.2) in Theorem 3.3) and some results of Fan et al. [13], we show that any convex DRM satisfying (1.3) admits a dual representation as in (1.6). Compared to [10], our result holds for unbounded terminal variables, and the function $f$ can be determined by the Legendre-Fenchel transformation of the generator $g$ in (3.2). Furthermore, for coherent DRMs satisfying (1.3), we obtain a dual representation as in (1.5), where the terminal variable $\xi$ is unbounded and $\mathcal{Q}_{1}$ is characterized by the subdifferential $\partial g(t,0)$. Our method relies on the representation (3.2) and therefore requires (1.3).
@@ 2 Quadratic $g$-expectations
${\cal{G}}^{\mu,\nu}_{\theta}:=\{g\in{\cal{G}}:$ $dt\times dP$-$a.e.,$ $\forall z,\tilde{z}\in\mathbf{R}^{d}$, $\forall\theta\in(0,1)$, $g(t,z)-\theta g(t,\tilde{z})\leqslant\mu|z-\theta\tilde{z}|+\frac{\nu}{1-\theta}|z-\theta\tilde{z}|^{2}\}$;
@@
$\leqslant\mu|z-\theta\tilde{z}|+\frac{\nu}{1-\theta}|z-\theta\tilde{z}|^{2}+(1-\theta)(\mu|\tilde{z}|+\nu|\tilde{z}|^{2})$
$\leqslant\mu(|z-\theta\tilde{z}|+|\tilde{z}-\theta z|)+\frac{\nu}{1-\theta}(|z-\theta\tilde{z}|^{2}+|\tilde{z}-\theta z|^{2})+(1-\theta)(\mu(|\tilde{z}|+|z|)+\nu(|\tilde{z}|^{2}+|z|^{2})).$
By interchanging $z$ and $\tilde{z}$, we have
$\lim_{\tilde{z}\rightarrow z}|g(t,z)-g(t,\tilde{z})|\leqslant 4(1-\theta)(\mu|z|+\nu|z|^{2})\rightarrow 0,\quad\text{as}\ \theta\rightarrow 1.$
@@
Let $Y_{t}$ be a $g^{\mu,\nu}$-submartingale. By Zheng et al. [39, Corollary 3.6], we have
$Y_{s}\leqslant\mathcal{E}^{\mu,\nu}[Y_{t}|{\cal{F}}_{s}]=\frac{1}{2\nu}\ln(E_{Q}[\exp(2\nu Y_{t})|{\cal{F}}_{s}]),\ \ \forall 0\leqslant s\leqslant t\leqslant T.$
@@
If $Y_{t}$ is a $g^{-\mu,-\nu}$-supermartingale, then $Y_{s}\leqslant-{\cal{E}}^{-\mu,-\nu}[Y_{t}|{\cal{F}}_{s}]={\cal{E}}^{\mu,\nu}[-Y_{t}|{\cal{F}}_{s}]$ for any $0\leqslant s\leqslant t\leqslant T$. This implies that $Y_{t}$ is a $g^{\mu,\nu}$-submartingale. Thus $Y_{t}$ admits an RCLL modification. ∎
@@
Since for all $s\in[0,t]$, $\exp(2\nu Y_{s})\leqslant\exp(2\nu\mathcal{E}^{\mu,\nu}[Y_{t}|{\cal{F}}_{s}])$, by (2.2), $\exp(2\nu Y_{t})$ is a $g^{\mu,0}$-submartingale. Then, by Peng [29, Theorem 3.3], there exists a pair $(z_{t},a_{t})\in H_{d}^{2}\times{\cal{A}}$, such that
@@
$Y_{t}=Y_{T}+$ $\int_{t+}^{T}\frac{1}{2\nu y_{s-}}\left(\mu|z_{s}|ds-da_{s}-z_{s}\cdot dB_{s}\right)+\int_{t+}^{T}\frac{|z_{s}|^{2}}{4\nu|y_{s}|^{2}}ds-\sum_{s\in(t,T]}\left(\Delta Y_{s}-\frac{1}{2\nu y_{s-}}\Delta y_{s}\right)$
@@ A DRM is called a regular DRM, if for all $\xi\in L^{0}(\mathcal{F}_{T})$, it satisfies (r7):
$\rho^{\mathrm{ent}}_{t}(\xi)=\mathcal{E}^{0,\nu}[-\xi|\mathcal{F}_{t}],\quad t\in[0,T],$
@@
In particular, when $\xi\in L^{\infty}(\mathcal{F}_{T})$, by (Proof.) in the Appendix, $\mathcal{Q}_{\xi,f}$ in (3.4) can be reduced to $\mathcal{Q}:=\{Q^{q}:$ $dQ^{q}/dP=\theta^{q}_{T}:=\int_{0}^{T}q_{s}\cdot dB_{s}-\frac{1}{2}\int_{0}^{T}|q_{s}|^{2}ds$ with $q_{t}\in\mathcal{H}_{d}^{BMO}\}.$ For $Q^{q}\in\mathcal{Q}$, the conditional relative entropy of $Q^{q}$ with respect to $P$:
@@ 5 A Doob-Meyer decomposition for $\rho$-supermartingales
$g(t,z)-\theta g(t,\tilde{z})\leqslant\mu|z-\theta\tilde{z}|+\frac{\nu}{1-\theta}|z-\theta\tilde{z}|^{2}.$
@@
Since $\rho$ satisfies (r5) and (r6) on $\mathcal{L}^{\exp}_{T}$, by Lemma 2.3(ix), we get that $g\in{\cal{G}}^{\mu,\nu}\cap{\cal{G}}_{subl}$. This implies that $dt\times dP$-$a.e.$, for all $z\in\mathbf{R}^{d}$ and $\beta>0$, $\beta|g(t,z)|=|g(t,\beta z)|\leqslant\beta\mu|z|+\beta^{2}\nu|z|^{2}$, and thus
$|g(t,z)|\leqslant\mu|z|+\beta\nu|z|^{2}.$
@@
where $B^{\tilde{q}^{n}}_{t}:=B_{t}-\int_{0}^{t}{\tilde{q}^{n}}_{s}ds$ is a standard Brownian motion under the probability $Q^{\tilde{q}^{n}}$ satisfying $Q^{\tilde{q}^{n}}/dP=\theta^{\tilde{q}^{n}}_{T}$. It follows from (Proof.) that there exists a constant $C$ independent of $n$ such that $E[\theta^{\tilde{q}^{n}}_{T}\ln(1+\theta^{\tilde{q}^{n}}_{T})]\leqslant C.$ Then by Protter [31, Theorem 11], we get that $\{\theta^{\tilde{q}^{n}}_{T}\}_{n\geqslant 1}$ is uniformly integrable, and then by Vitali convergence theorem, we have $E[\theta^{\tilde{q}}_{T}]=1$, which implies that $\theta^{\tilde{q}}_{t}$ is a uniformly integrable martingale. Moreover, by Fatou’s lemma, we also have
@@
By (Proof.), (a.10), Fatou’s lemma and the fact that $f\geqslant 0$, we have
@@
From (a.10) and (a.11), it follows that $\tilde{q}_{t}\in\mathcal{Q}_{(\xi,f)}$. Then, by (Proof.)-(a.11) and domination convergence theorem, we have
@@
Since $f(t,q_{t})\geqslant q_{t}\cdot Z_{t}-g(t,Z_{t}),$ by (Proof.) and similar arguments as in (Proof.)-(a.10), we can obtain (a.13), which together with (a.12) and Definition 2.2(i) gives (6.9). ∎